### v1

@@ 3.1  Data and measures
Table 2 presents the sample size and the distribution of players by color according to the FairFace algorithm (Panel A) and human raters (Panel B) for the entire period (2004-2014). We provide separate tables of sample size for each investigated period (2004-2006, 2007-2010, and 2011-2014) in Appendix Tables B.2, B.3 and B.4. In total, our sample includes 505 players and 69 referees who participated in 2,359 games over 11 seasons.
According to the FairFace algorithm, there are 335 Black and 170 non-Black players, as well as 23 Black and 46 non-Black referees. According to human raters, the numbers are similar, but not identical (335 Black and 170 non-Black players, as well as 27 Black and 42 non-Black referees). By looking at the distribution of L* values, for Black and non-Black individuals, we see upper and lower bounds (min-max). Importantly, however, there is a considerable overlap of L* between Black and non-Black individuals, which justifies our focus on skin tone. Since race is subjective [61], reducing the analysis to just race oversimplifies the task and might even introduce bias, given the large amount of evidence about inter and intra-race skin tone preferences [22, 47, 64]. Finally, we also created a variable that represents the average absolute difference in skin tone (L*) between a player and each member of the three-person referee crew. As described in Panel C of Table 2, there were 1,876 unique crews with the mean absolute distance of 8.39 (sd = 4.19).
Table 3 presents the summary statistics for the different players, referees and games characteristics along with a comparison of the mean values between Black and non-Black players. These summary statistics reveal significant differences between Black and non-Black players in minutes played, number of fouls, points, and different adjusted per 40 minutes characteristics that relate to players’ performance. We will return to these differences when discussing our empirical strategy. We also see differences in the referee crew composition for Black and non-Black players. In Appendix Tables B.5, B.6, B.7, B.8, B.9, B.10 and B.11, we present summary statistics for each investigated period (2004-2006, 2007-2010, and 2011-2014) per each race classification (FairFace and human raters). In general, the results presented in the tables for each investigated period, show no significant pattern of associations between the race of players and the referee crew.
@@ 3.2  Empirical strategy
Studying the effect of the difference between a player and a referee crew’s race/color combination on the number of fouls is challenging. A naïve approach of correlating the number of fouls and the difference between a player and a referee crew’s race/color would yield biased and inconsistent estimates. This is because players’ characteristics (e.g., defensive and offensive abilities, position, superstar status, experience, etc.) vary between Black and non-Black players (see Table 3). Moreover, the individual ability and the status of players within a team may vary over time. [^21] See [32] for the discussion on change in ability of professional athletes throughout their careers and the use of fixed effects over long versus short periods. Therefore, different sources of unobserved heterogeneity need to be taken into account.
@@
In addition, it is important to refer to the referee selection process. The advantages of our setting lie in (i) the process of assigning referees to games, which does not consider player race/skin tone (see Table 1), thereby ensuring that our findings are not confounded by subjects sorting to preferred evaluators or vice versa, and (ii) repeated interactions bypassing selection issues. Accordingly, the most basic specification of our model takes the following form:
$\text{Foul rate}_{itgy}=\alpha\text{Black player}_{i}\times\%\text{non-Black referees}_{g}+\theta_{iy}+\gamma_{g}+X_{igt}+\epsilon_{itgy},$ (1)
where the dependent variable is the number of fouls earned per 40 minutes (40×fouls/minutes played) by player $i$ from team $t$, in a specific game $g$, in year $y$. The coefficient of interest, $\alpha$, is for the interaction variable between an indicator of whether player $i$ is Black and the fraction of non-Black referees in a specific game $g$. $\theta_{iy}$ denotes player-year fixed effects, controlling for time-invariant player characteristics that are constant within a year (season), and $\gamma_{g}$ denotes game fixed effects, controlling for game-invariant characteristics (e.g., referee crew, attendance, and team performance up until the specific game). $X_{igt}$ is our set of controls, which includes a dummy variable indicating whether a player is a starter, whether her team played the game at home, and the race of her coach. [^22] While [73] showed a positive relationship between a player’s playing time and the similarity between the race of the player and the coach in the NBA, [5] found no such relationship in the WNBA for the 2018-2022 seasons. All estimates are weighted by the number of minutes played, and standard errors are clustered at the player and game level. [^23] In contrast to [55], we opt for two-way clustering at the player and game levels because of possible dependency between the behavior of players and referee crews within a game [1].
@@
$\text{Foul rate}_{itgy}=\beta\;\text{CrewDistance}_{ig}+\theta_{iy}+\gamma_{g}+X_{igt}+\epsilon_{itgy},$ (2)
Based on [54], we utilize a natural experiment arising from the widespread publicity of [55], which served as an external shock intensifying scrutiny around racial bias in men’s basketball officiating. This event potentially disrupted established norms of referee discrimination in the NBA, prompting an investigation into whether similar changes occurred in the women’s context. We divide our analysis into three periods: a baseline period (2004-2006) before media exposure, a treated period (2007-2010) immediately following increased public awareness, and an extended period (2011–2014) to examine the persistence or evolution of any observed effects. We combine the relevant periods to run a pooled regression, extending equations (1) and (2) by adding an interaction term between a relevant subsequent period (i.e., 2004-2006 along with 2007-2010 or 2007-2010 along with 2011-2014) and our variable of interest $\text{Black player}_{i}\times\%\text{non-Black referees}_{g}$, presented in equation (1).
@@ 4.1  Binary race measure
We start our analysis using the binary race classification of Black or non-Black as determined by the FairFace algorithm. Column 1 of Table 4 shows that in the period before media exposure (pre-awareness period between 2004 and 2006), there was no statistically significant evidence of racial discrimination in WNBA foul calling ($\alpha$ = 0.076, p = 0.713). While this result is in contrast to [55], it could be an expected result given the organization of interest is the WNBA, an organization with longstanding commitment to equality. In Column 2, we see a negative coefficient of $\alpha$, which suggests overcorrection in the immediate period following media exposure (2007-2010), according to which players are called for fewer fouls, the higher the share of referees with unlike color. However, this estimate is not significant at conventional levels (p = 0.205). Column 4 shows that the coefficient of $\alpha$ in the 2007-2010 period is lower than in the 2004-2006 period. But here again, the change is not significant at conventional levels (p = 0.325). Columns 3 and 5 represent the results for the extended period following media exposure (2011-2014). Overall, these estimates are also not precise enough to become significant at conventional levels, but the general trends are consistent throughout. Moreover, when pooling the data from 2007 to 2014, the results in Column 5 are significant at the 10% level, suggesting a higher coefficient of $\alpha$ in 2011-2014 compared to 2007-2010.
@@
In Table 5, we present similar analyses using race categorizations from human raters. We again find that $\alpha$ is not significant in the pre-awareness period ($\alpha$ = 0.106, p = 0.541, see Column 1). However, unlike the results with FairFace algorithm, $\alpha$ becomes negative and statistically significant in the 2007–2010 post-awareness period ($\alpha$ = -0.258, p = 0.046, see Column 2). The results of the pooled regression that combines data from 2004 to 2010 suggest that there was a significant reduction in the foul calls against players that were refereed by opposite color crews between pre- and post-awareness periods (Post Black × fraction non-Black referees = -0.365, p = 0.098, see Column 4). These results serve as suggestive evidence of overreaction by referees that started calling significantly fewer fouls against players of a different color in the period after the widespread media attention. However, we observe a significant reversal of this overreaction in the 2011–2014 period, which is precise enough to become significant in the pooled regression (Post Black × fraction white referees = 0.474, p = 0.011, see Column 5) suggesting that the effect wore off over time. Note that when looking at the 2011-2014 period separately, while $\alpha$ is positive, it is not significant at conventional levels ($\alpha$ = 0.218, p = 0.136, see Column 3).
@@
  • Notes: See comments for Table 4. ***p <0.01; **p <0.05; *p <0.1.
Given the observed differences in racial categorizations between the FairFace algorithm and the human raters in Table 2, we further analyze a sample including only observations where both agreed on classifications. This approach addresses the inherent challenge of assigning individuals to a single racial category. Table 6 presents the results of these estimations for this new, slightly reduced, dataset. We find that in the pre-awareness period, presented in Column 1, $\alpha$ is positive (0.236) but not significant (p = 0.250), similar in direction to earlier specifications. However, immediately after media attention, the coefficient turns significantly negative ($\alpha$ = –0.264, p = 0.063, see Column 2), indicating that, for games with a higher share of the same color referees, players from a different color were less likely to be called for fouls. In the extended post-awareness window, the coefficient again reverses positive ($\alpha$ = 0.334, p = 0.056, see Column 3), suggesting a rebound in foul-calling patterns. The pooled regressions further underscore these shifts (presented in Columns 4-5), which are statistically meaningful and exceed the magnitudes observed in either the FairFace-only (Table 4) or human-rater-only (Table 5) samples.
Table 6:
Pre-awareness Post-awareness Change in coefficient
Black $\times$ 0.236 -0.264* 0.334* 0.237 -0.262*
fraction non-Black referees (0.204) (0.141) (0.174) (0.204) (0.141)
Post Black $\times$ -0.499* 0.593***
fraction non-Black referees (0.257) (0.211)
N 11,425 16,508 12,861 27,933 29,369
Sample mean 4.38 4.52 4.10 4.46 4.34
  • Notes: See comments for Table 4. ***p <0.01; **p <0.05; *p <0.1.
One possible issue could be with the definitions of the periods. For example, [54] used a three-year pre-awareness period and a four-year post-awareness period. To ensure this uneven period length did not drive our results, we re-evaluated foul call trends by dividing the years following awareness into consecutive three-year segments: 2004-2006, 2007-2009, and 2010-2012. Our analysis, presented in Appendix Tables C.1 and C.2, demonstrates that, overall, our results hold consistently regardless of whether a 3-year or 4-year post-awareness evaluation period is used.
@@ 4.2  Continuous skin tone measure
In Table 7, we present the results for equation (2) using a continuous skin tone measure. As for the binary race measures, our results using skin tone for the pre-awareness period are highly insignificant ($\beta$ = -0.000, p = 0.963; see column 1). However, in Column 2, we observe a statistically significant negative coefficient that emerges in the immediate post-awareness period ($\beta$ = -0.020, p = 0.003), suggesting that there is evidence of overreaction immediately after widespread media attention, also when using a continuous skin tone measure. To put this result into perspective, a one standard deviation change in the difference between referees and players L* (sd = 4.27) corresponds to 0.085 fewer fouls per 40 minutes played. In Column 4, we also see that there is a significant difference in $\beta$ in the 2007-2010 period compared to the 2004-2006 period (Post crew distance = -0.019, p = 0.030). Finally, the extension to 2014, presented in Column 5, shows a significant increase in $\beta$ in the 2011-2014 period compared to the 2007-2010 period (Post crew distance = 0.026, p = 0.009) back to near pre-awareness levels, as seen in Column 3. Overall, these findings mimic our findings using the binary race categories, though the estimates using the continuous skin tone measure are comparably more precise to become significant even at the 1-percent level.
@@
  • Notes: Dependent variable is defined as 40×fouls/minutes. Crew distance is the mean of the absolute distance in skin tone values between the player and each referee in the officiating crew. Each regression includes player-year fixed effects, game fixed effects, whether a player is a starter, whether a player is playing in her home arena, and her coach’s skin tone. The last two columns report the results of a pooled regression, including both a primary period and a subsequent period to measure the changes throughout our main sample years. Each observation is weighted by the number of minutes played. Standard errors are clustered at the player and game levels and appear in parentheses. ***p <0.01; **p <0.05; *p <0.1.
@@ 4.3  In-group discrimination versus out-group favoritism
An often-discussed issue in empirical studies of preferences and bias is the inability to disentangle in-group versus out-group bias [26, 58]. The net effect may be the same, but the mechanism is different. For example, [26] distinguish between “exophilic” case, where agents prefer members of other groups, and “endophobic” case, where agents discriminate against people like themselves. In our setting, we are able, at least to some extent, to differentiate between in- and out-group bias by adding the quadratic transformation of the CrewDistance variable in equation (2). The intuition of this approach is as follows: by exploring any non-linearity between foul rate and CrewDistance, we can identify whether the effect occurs at small(er) distances, involving in-group members (i.e., crews and players of similar skin tone), or at large(r) distances, involving out-group members (i.e., crews and players of different skin tone), or both.
Figure 5 illustrates the non-linearity between foul rate and absolute crew mean distance for the pre-awareness 2004-2006 period (left column), the immediate post-awareness 2007-2010 period (middle column), and the rest of the sample period covering 2011-2014 seasons (right column). The top row shows a locally weighted smoothing (LOWESS) of the relationship between foul rate (y-axis) and the mean of absolute distance in skin tone values between the player and each referee in the officiating crew (x-axis) for each sample, illustrating the raw nonparametric pattern. The bottom row depicts the marginal effect of distance derived from the quadratic model, which follows the specification from equation (2), along with 95% confidence intervals. Overall, these plots reveal how the average skin tint difference between each referee and the player influences changes in foul rates across its spectrum. While defining a clear-cut regarding “group affiliation” seems arbitrary, in fact, one would expect that such a cut-off is highly idiosyncratic and varies between individuals, it seems reasonable to assume that any crew-athlete constellation closer to zero represents a skin tint-related in-group while a crew-athlete constellation at larger distances represents a skin tint-related out-group.
@@
The first two panels exhibit a downward slope as distance increases, indicating that distances (beyond certain values) lead to a negative marginal effect, i.e., an increase in the difference between referees and players’ tone color correlates with a decrease in predicted foul rate. Near smaller distances (0–5), the marginal effect hovers around or slightly above zero, suggesting that when distance is minimal, slight differences in skin tone between players and referees do not systematically affect foul calls. If any, there is a slight tendency to punish players with similar skin tones in the pre-awareness period, though the results are not significant. As distance increases (particularly past 10 or 15, depending on the panel), the marginal effect drops below zero, indicating a negative impact on the foul rate per unit of distance, which is also statistically significant in the 2007-2010 period, suggesting an out-group favoritism rather than in-group discrimination. We see no clear and significant pattern in the third panel that represents the 2011-2014 period.

### v3

@@ 3.1  Data and measures
In total, our sample includes 505 players and 69 referees who participated in 2,359 games over 11 seasons. Table 2 presents the sample size and the distribution of players by color according to the FairFace algorithm (Panel A) and human raters (Panel B) for the entire period (2004-2014). We provide separate tables of sample size for each investigated period (2004-2006, 2007-2010, and 2011-2014) in Appendix Tables B.2, B.3 and B.4.
According to the FairFace algorithm, there are 335 Black and 170 non-Black players, as well as 23 Black and 46 non-Black referees. According to human raters, the numbers are similar, but not identical (335 Black and 170 non-Black players, as well as 27 Black and 42 non-Black referees). By looking at the distribution of perceptual skin tint values for Black and non-Black individuals, we see upper and lower bounds (min-max) as well as a considerable overlap between Black and non-Black individuals. Finally, we created a variable that represents the average absolute difference in skin tone (L*) between a player and each member of the three-person referee crew. As described in Panel C of Table 2, there were 1,876 unique crews with the mean absolute distance of 8.39 (sd = 4.19).
Table 3 presents the summary statistics for the different players, referees, and game characteristics along with a comparison of the mean values between Black and non-Black players. These summary statistics reveal significant differences between Black and non-Black players in minutes played, number of fouls, points, and different adjusted per 40 minutes characteristics that relate to players’ performance. We will return to these differences when discussing our empirical strategy. We also see differences in the referee crew composition for Black and non-Black players. In Appendix Tables B.5, B.6, B.7, B.8, B.9, B.10 and B.11, we present summary statistics for each investigated period (2004-2006, 2007-2010, and 2011-2014) per each race classification (FairFace and human raters). In general, the results presented in the tables for each investigated period show no significant pattern of associations between the race of players and the referee crew.
@@ 3.2  Empirical strategy
Studying the effect of the difference between a player and a referee crew’s race or color combination on the number of fouls is challenging. A naïve approach of correlating the number of fouls and the difference between a player and a referee crew’s race or color would yield biased and inconsistent estimates. This is because players’ characteristics (e.g., defensive and offensive abilities, position, superstar status, experience, etc.) vary between Black and non-Black players (see Table 3). Moreover, the individual ability and the status of players within a team may vary over time. [^20] See [gschwend2021] for the discussion on change in ability of professional athletes throughout their careers and the use of fixed effects over long versus short periods. As such, for identifying the effects of interest, different sources of unobserved heterogeneity need to be taken into account.
@@
In addition, it is important to refer to the referee selection process. The advantages of our setting lie in (i) the process of assigning referees to games, which does not consider player race or skin tone (see Table 1), thereby ensuring that our findings are not confounded by subjects sorting to preferred evaluators or vice versa, and (ii) repeated interactions bypassing selection issues. Accordingly, the most basic specification of our model takes the following form:
$\text{Foul rate}_{itgy}=\alpha\text{Black player}_{i}\times\%\text{non-Black referees}_{g}+\theta_{iy}+\gamma_{g}+\textbf{$\rho^{\prime}$}\textbf{X}_{igt}+\epsilon_{itgy},$ (1)
where the dependent variable is the number of fouls earned per 40 minutes (40×fouls/minutes played) by player $i$ from team $t$, in a specific game $g$, in year $y$. The coefficient of interest, $\alpha$, is for the interaction variable between an indicator of whether player $i$ is Black and the fraction of non-Black referees in a specific game $g$. We interpret it as the differential impact of the racial composition of the refereeing crew on Black players relative to non-Black players. The coefficient is identified from the variation in referee crew composition between games.
$\theta_{iy}$ denotes player-year fixed effects, controlling for time-invariant player characteristics that are constant within a year (season), and $\gamma_{g}$ denotes game fixed effects, controlling for game-invariant characteristics (e.g., referee crew, attendance, and teams’ performance up until the specific game). $\textbf{X}_{igt}$ is our vector of controls, which includes a dummy variable indicating whether a player is a starter, whether her team played the game at home, and the race of her coach. [^21] While [zhang2017] showed a positive relationship between a player’s playing time and the similarity between the race of the player and the coach in the NBA, [agha2024] found no such relationship in the WNBA for the 2018-2022 seasons. All estimates are weighted by the number of minutes played, and standard errors are clustered at the player and game level. [^22] In contrast to [price2010], we opt for two-way clustering at the player and game levels because of possible dependency between the behavior of players and referee crews within a game [abadie2023].
@@
$\text{Foul rate}_{itgy}=\beta\;\text{CrewDistance}_{ig}+\theta_{iy}+\gamma_{g}+\textbf{$\rho^{\prime}$}\textbf{X}_{igt}+\epsilon_{itgy},$ (2)
Based on [ppw2018], we utilize a natural experiment arising from the widespread publicity of [price2010], which served as an external shock intensifying scrutiny around racial bias in men’s basketball officiating. This event potentially disrupted established norms of referee discrimination in the NBA, prompting an investigation into whether similar changes occurred in the women’s context. We divide our analysis into three periods: a baseline period (2004-2006) before media exposure, a treated period (2007-2010) immediately following increased public awareness, and an extended period (2011–2014) to examine the persistence or evolution of any observed effects. We combine the relevant periods to run a pooled regression, extending equations (1) and (2) by adding an interaction term between a relevant subsequent period (i.e., 2004-2006 along with 2007-2010 or 2007-2010 along with 2011-2014) and our variables of interest $\text{Black player}_{i}\times\%\text{non-Black referees}_{g}$ and $\text{CrewDistance}_{ig}$, respectively.
@@ 4.1  Binary race measure
We start our analysis using the binary race classification of Black or non-Black as determined by the FairFace algorithm. Recall, a positive (negative) sign of $\hat{\alpha}$ indicates that players are called for more (fewer) fouls, the higher the share of referees with unlike race classifications. As such, and in line with [price2010] a positive (negative) estimate would be suggestive of racial discrimination (favoritism) in foul calling. In their main specification, [price2010] report a significant positive estimate of 0.181 (p << 0.05) (see Table 4 in their paper).
As can be seen in Column 1 of Table 4, we do not find any significant evidence of racial discrimination in WNBA foul calling for the period before media exposure (pre-awareness period between 2004 and 2006). Compared with the NBA estimate for the same period, our WNBA estimate is small and not significant ($\hat{\alpha}$ = 0.077, p = 0.713). Interestingly, however, in Column 2, we see a negative estimate of -0.183 (p = 0.205) for the immediate period following media exposure (2007-2010). Notably, even though not significant at conventional levels, this estimate is much larger in absolute value and more precise compared to the pre-awareness period. Content-wise, this is in line with overcorrection as discussed before. Finally, in Column 3, we see again a positive but not significant estimate for the extended period following media exposure (2011-2014) ($\hat{\alpha}$ = 0.212, p = 0.199). While not significant at conventional levels, this result suggests an overreaction in the opposite direction. We will come back to this result in the next sub-section when estimating the continuous skin tone measure to see whether this result is robust or not.
@@
Overall, when analyzing the pooled data combining the periods between 2004 and 2010 as well as between 2007 and 2014, these temporal patterns remain: $\hat{\alpha}$ is lower in the 2007-2010 period than in the 2004-2006 period (see Column 4) and in the 2011-2014 period (see Column 5). The latter finding is significant at the 10%-level.
Using race categorizations from human raters instead of FairFace, we confirm the general patterns from before (see Table 5). However, while $\hat{\alpha}$ is still neither significant in the pre-awareness period ($\hat{\alpha}$ = 0.106, p = 0.541, see Column 1 ) nor in the extended period following media exposure (i.e., 2011-2014 with $\hat{\alpha}$ = 0.218, p = 0.136, see Column 3), the estimate in the 2007–2010 post-awareness period is precise enough to become significant at conventional levels ($\hat{\alpha}$ = -0.258, p = 0.046, see Column 2). Again, when analyzing the pooled data combining the periods between 2004 and 2010 as well as between 2007 and 2014, these temporal patterns remain: $\hat{\alpha}$ is lower in the 2007-2010 period than in the 2004-2006 period (see Column 4) and in the 2011-2014 period (see Column 5). These findings are significant at 10% and 5% levels, respectively.
@@
  • Notes: Dependent variable is defined as 40×fouls/minutes. Each regression includes player-year fixed effects, game fixed effects, whether a player is a starter, whether a player is playing in her home arena, and the skin color of her coach. The last two columns report the results of a pooled regression, including both a relevant primary period and a subsequent period to measure the change in our variable of interest between the two periods. Each observation is weighted by the number of minutes played. Standard errors are clustered at the player and game levels and appear in parentheses.
    ***p <0.01; **p <0.05; *p <0.1.
Using a reduced dataset with a sample that only includes observations where both FairFace and human raters agreed on classifications, we find the same temporal patterns as before but (slightly) increased effect sizes (see Table C.1 for more details). Moreover, we test whether referee decisions depend on collective appearance, i.e., a player’s race relative to the racial mix of both competing teams. Our findings hold even after controlling for the racial gap between competing teams (see Appendix Tables C.2 and C.3). Another possible issue could be with the definitions of the periods. For example, [ppw2018] used a three-year pre-awareness period and a four-year post-awareness period. To ensure this uneven period length did not drive our results, we re-evaluated foul call trends by dividing the years following awareness into consecutive three-year segments: 2004-2006, 2007-2009, and 2010-2012. Our analysis, presented in Appendix Tables C.4 and C.5, demonstrates that, overall, our results hold consistently regardless of whether a 3-year or 4-year post-awareness evaluation period is used.
Summing up, even though several estimates are not precise enough to become significant at conventional levels, the temporal patterns and general trends are consistent throughout: while there is no statistically significant evidence of racial discrimination in WNBA foul calling for the period before media exposure, we do observe a tendency for overcorrection in the period following media exposure, as well as a rebound effect in foul-calling patterns for the extended period following media exposure.
@@ 4.2  Continuous skin tone measure
In Table 6, we present the results for equation (2) using a continuous skin tone measure. As for the binary race measures, our results using skin tone for the pre-awareness period indicate that no discrimination exists ($\hat{\beta}$ = -0.000, p = 0.963; see column 1). However, in Column 2, we observe a statistically significant negative coefficient that emerges in the immediate post-awareness period ($\hat{\beta}$ = -0.020, p = 0.003), suggesting that there is evidence of overcorrection immediately after widespread media attention, also when using a continuous skin tone measure.
@@
  • Notes: Dependent variable is defined as 40×fouls/minutes. Crew distance is the mean of the absolute distance in skin tone values between the player and each referee in the officiating crew. Each regression includes player-year fixed effects, game fixed effects, whether a player is a starter, whether a player is playing in her home arena, and her coach’s skin tone. The last two columns report the results of a pooled regression, including both a primary period and a subsequent period to measure the changes throughout our main sample years. Each observation is weighted by the number of minutes played. Standard errors are clustered at the player and game levels and appear in parentheses.
    ***p <0.01; **p <0.05; *p <0.1.
To put this result into perspective, a one standard deviation change in the difference between referees and players L* (sd = 4.27) corresponds to 0.085 fewer fouls per 40 minutes played. In Column 4, we also see that there is a significant difference in $\hat{\beta}$ in the 2007-2010 period compared to the 2004-2006 period (Post crew distance = -0.019, p = 0.030). Finally, the extension to 2014, presented in Column 5, shows a significant increase in $\hat{\beta}$ in the 2011-2014 period compared to the 2007-2010 period (Post crew distance = 0.026, p = 0.009) back to near pre-awareness levels, as seen in Column 3. Overall, these findings mimic our findings using the binary race categories, though the estimates using the continuous skin tone measure are comparably more precise to become significant even at the 1-percent level. [^23] In addition, we explicitly test for potential team-level bias in racial composition by adding team-year fixed effects into our estimations. This accounts for the varying share of black players on rosters and minutes played for a specific team in a given year. Including team-year fixed effects did not alter our results. The results of these estimations are available upon request. This holds even when accounting for the skin tone composition of the teams (Appendix Table C.7, Panel B). Finally, it is important to note that the pattern of overcorrection in the 2007-2010 period remains the same in estimations with binary race and continuous skin tone measures. However, the null effect in the 2011-2014 period (Column 3 in Table 6) suggests that the large overcorrection in the opposite direction, as was observed with the binary race measure, is not replicable with the continuous skin tone measure.
@@ 4.3  In-group discrimination versus out-group favoritism
An often-discussed issue in empirical studies of preferences and bias is the inability to disentangle in-group versus out-group bias [feld2016, sandberg2018]. The resultant net effect may be the same, but the mechanism is different. For example, [feld2016] distinguish between “exophilic” case, where agents prefer members of other groups, and “endophobic” case, where agents discriminate against people like themselves. In our setting, we are able, at least to some extent, to differentiate between in- and out-group bias by adding the quadratic transformation of the CrewDistance variable in equation (2). The intuition of this approach is as follows: by exploring any non-linearity between foul rate and CrewDistance, we can identify whether the effect occurs at small(er) distances, involving in-group members (i.e., crews and players of similar skin tone), or at large(r) distances, involving out-group members (i.e., crews and players of different skin tone), or both.
@@
Figure 5 illustrates the non-linearity between foul rate and absolute crew mean distance for the pre-awareness 2004-2006 period (left column), the immediate post-awareness 2007-2010 period (middle column), and the rest of the sample period covering 2011-2014 seasons (right column). The top row shows a locally weighted smoothing (LOWESS) of the relationship between foul rate (y-axis) and the mean of absolute distance in skin tone values between the player and each referee in the officiating crew (x-axis) for each sample, illustrating the raw nonparametric pattern. The bottom row depicts the marginal effect of distance derived from the quadratic model, which follows the specification from equation (2), along with 95% confidence intervals. Overall, these plots reveal how the average skin tint difference between each referee and the player influences changes in foul rates across the spectrum. While defining a clear-cut regarding “group affiliation” seems arbitrary, in fact, one would expect that such a cut-off is highly idiosyncratic and varies between individuals, it seems reasonable to assume that any crew-athlete constellation closer to zero represents a skin-tint-related in-group while a crew-athlete constellation at larger distances represents a skin-tint-related out-group.
The first two panels exhibit a downward slope as distance increases, indicating that distances (beyond certain values) lead to a negative marginal effect, i.e., an increase in the difference between referees’ and players’ tone color correlates with a decrease in predicted foul rate. Near smaller distances (0–5), the marginal effect hovers around or slightly above zero, suggesting that when distance is minimal, slight differences in skin tone between players and referees do not systematically affect foul calls. If any, there is a slight tendency to punish players with similar skin tones in the pre-awareness period, though the results are not significant. As distance increases (particularly past 10 or 15, depending on the panel), the marginal effect drops below zero, indicating a negative impact on the foul rate per unit of distance, which is also statistically significant in the 2007-2010 period, suggesting an out-group favoritism rather than in-group discrimination. We see no clear and significant pattern in the third panel that represents the 2011-2014 period.